.Consider the following: Net operating income under variable costing $50,000. Decrease in inventory during the period 5,000 units. Fixed manufacturing overhead $100,000. Number of units produced during the period 25,000 units. Based on the above information, the net income under absorption costing is: a. $150,000 b. $70,000 c. $30,000 d. $50,000
Answer #1

Answer: Option ( B ) $70,000

Net Income under absorption costing = Net operating income under variable costing + Fixed manufacturing overhead

Net Income under absorption costing = $50,000 + $20,000 = $70,000

Fixed manufacturing overhead = (100000 / 25000) = $4 * 5000 = $20,000